Adam is twice as old as his brother. The sum of their ages is 27. How old is each boy?

Answer by checkley79(3341)

A=2X
B=X
2X+X=27
3X=27
X=27/3
X=9 YEARS OLD FOR THE BROTHER.
2*9=18 YEARS OLD FOR ADAM.